OpenStudy (anonymous):

1. What is the value of g(-3) when g(x)=2x-2? g(-3)= ?? 2. Evaluate the function g(t)=4/5t when t=50 A. 45 << My answer B. 40 C. 35 D. 30 3. Evaluate the function H(x)=16-x when x=-4 A. 64 B. 20 C. 12 D. -64

OpenStudy (ipwnbunnies):

Just plug in -3 for x. g(-3) = 2*-3 + 2? Idr if it was plus 2 or minus 2. All the characters are messed up >.<

OpenStudy (anonymous):

Yeah, sorry all of the characters just messed up as I was editing. The questions (and answer choices are): 1. What is the value of g(-3) when g(x)=2x-2? g(-3)= ?? 2. Evaluate the function g(t)=4/5t when t=50 A. 45 << My answer B. 40 C. 35 D. 30 3. Evaluate the function H(x)=16-x when x=-4 A. 64 B. 20 C. 12 D. -64

OpenStudy (ipwnbunnies):

So, yeah, just plug in -3 for x. 2*-3 - 2 = ? For the second question, plug in 50 for t. Last question, same deal. plug in -4 for x. Be careful though, it'll be 16-(-4), which becomes 16 + 4. A negative times a negative equals a positive.

OpenStudy (anonymous):

So what would number two be? Is choice A (45) correct?

OpenStudy (anonymous):

Also for number three, I get H=-5.